OpenStudy (anonymous):

find the solution set by factoring and using zero product rule or by using the quadratic formula 6x^2+5x-4=0

OpenStudy (anonymous):

x= -4/3 y = 1/2

OpenStudy (anonymous):

shud i show sol

OpenStudy (anonymous):

i have to enter my answer in fractional form using exact roots

OpenStudy (anonymous):

these are exact and in fractional form